OpenStudy (bossimbacon):

Anna’s Work: 8 ÷ -2/3 => 8/1 • -2/3 => -16/3a) Find the error that Anna made when simplifying. @jim_thompson5910 @mathstudent55

jimthompson5910 (jim_thompson5910):

she forgot to flip the second fraction -2/3 to get -3/2

jimthompson5910 (jim_thompson5910):

when you divide two fractions, you flip the second fraction and multiply \[\Large \frac{A}{B} \div \frac{C}{D} = \frac{A}{B} \times \frac{D}{C}\]

OpenStudy (bossimbacon):

ahhh I see thank you!

jimthompson5910 (jim_thompson5910):

np